a printer has ink that is 5% cobalt blue color and ink that is 25% cobalt blue color. how many ounces of each ink are needed to make 3 gallons (384 ounces) of ink that is 15% cobalt blue color?
Answer by ankor@dixie-net.com(22740)

a printer has ink that is 5% cobalt blue color and ink that is 25% cobalt blue color.
how many ounces of each ink are needed to make 3 gallons (384 ounces) of ink that is 15% cobalt blue color?
:
let x = amt of 25% required
the total is to be 384 oz, therefore
(384-x) = amt of %5
:
.25x + .05(384 - x) = .15(384)
.25x + 19.2 - .05x = 57.6
.25x - .05x = 57.6 - 19.2
.2x = 38.4
x = 38.4/.2
x = 192 oz of the 25% stuff
then
384 - 192 = 192 oz of the 5% also